OpenStudy (anonymous):

A bicyclist of mass 90 kg drives around a circle with a centripetal acceleration of 1.5 m/s2. What is the centripetal force acting on the bicyclist? -with answer please so i can plug it in and it could help me understand better-

OpenStudy (michele_laino):

the centripetal force F_c is given by the subsequent formula: \[\Large {F_c} = m \times {a_c}\] where a_c is the centripetal acceleration, and m is the mass of the bicyclist

OpenStudy (michele_laino):

\[\Large {F_c} = m \times {a_c} = 90 \times 1.5 = ...Newtons\]